OpenStudy (anonymous):

***RECURSIVE EQUATION*** A certain business has implemented a new incentive plan whereby an employee earns tokens for every sale they make. For 1 sale, an employee earns 1 token. For 2 sales, an employee earns 6 tokens. For 3 sales, an employee earns 12 tokens. For 4 sales, an employee earns 19 tokens. Which recursive equation represents the pattern? an = an - 1 + 3n + 1 an = an - 1 + 3(n + 1) an = an - 1 + (n + 3) an = an - 1 + (n – 3)

OpenStudy (amistre64):

add 5, add6, add7

OpenStudy (amistre64):

a2 = 1 + 5 a3 = 1 + 5 + 6 a4 = 1 + 5 + 6 + 7

OpenStudy (amistre64):

an = an-1 + (n + 3) maybe a2 = 1 + (5) a3 = 6 + (6) a4 = 12 + (7) yeah